OpenStudy (anonymous):

8^x =2/16^(1-x) help!

OpenStudy (anonymous):

try writing the equation out like this: 16^(1-x) * 8^x = 2 then change that the 16 and 8 to a base 2. power of two... not base 2...

OpenStudy (anonymous):

any progress?

OpenStudy (anonymous):

16^(1-x) * 8^x=2 2^4-4x *2^3x=1 4-4x+3x=1 x=3 ??

OpenStudy (jlastino):

Yeah